K FV gN^{∗1}

wow^{∗2}

1

gNXAjreportAsectiongB pbP[WAamsmath.styAamsthm.styAamssymb.stygeometry.stygB RhhAgeometry.stygB

2 Alto

ObtBNEC[tF CXAltoB1973NAAEP CCiubNRZvgA[bNXoEAgubCiubNvB

f(x) = ddd
f(x) = ddd  (1)

ObtBNEC[tF CXAltoB1973NAAEP CCiubNRZvgA[bNXoEAgubCiubNvB

3 test

AEX—CeBOEfoCXA RhIwB AXN[An‘tHgA AWYSIWYGvVIB

fAgAOXEGQo[gT wL[[hMdB TEXALamport (1994)Knuth (1984)

4 Lisa

1979NAAbvXeB[uEWuXAltoG LisaB1983N1LisaLocally Integrated Software Architecture AWuXOB LisaAEXRhAArbgbv…

5 Alto

ObtBNEC[tF CXAltoB1973NAAEP CCiubNRZvgA[bNXoEAgubCiubNvB

6 test

AEX—CeBOEfoCXA RhIwB AXN[An‘tHgA AWYSIWYGvVIB

fAgAOXEGQo[gT wL[[hMdB TEXALamport (1994)Knuth (1984)

7 Lisa

1979NAAbvXeB[uEWuXAltoG LisaB1983N1LisaLocally Integrated Software Architecture AWuXOB LisaAEXRhAArbgbv…

8 Ql

Leslie Lamport. L^{A}TEX: A Document Preparation System. Addison-Wesley, Reading, Massachusetts, second edition, 1994, ISBN 0-201-52983-1.

Donald E. Knuth. The TEXbook, Volume A of Computers and Typesetting, Addison-Wesley, Reading, Massachusetts, second edition, 1984, ISBN 0-201-13448-9.

∗1 Yoshiki ”Newbie” KUMAZAWA ∗2 w

Yoshiki Kumazawa 1-1-1 Bamba, Hikone Faculty of Economics, Shiga University Shiga, JAPAN E-mail: kumazawa@biwako.shiga-u.ac.jp